OpenStudy (lovelyanna):

Which ordered pair is on the graph of the equation 2x + 5y = 4? A. (0, –1) B. (0, 1) C. (7, –2) D. (3, –2)

OpenStudy (lovelyanna):

Good morning Jiggly!

OpenStudy (lovelyanna):

Or afternoon..

jigglypuff314 (jigglypuff314):

Hello to you too ;) to figure which ordered pair is on the graph, you plug in the ordered pair (x, y) so for example for (0, -1) 2x + 5y = 4 --> 2(0) + 5(-1) = 4 0 + -5 = 4 -5 does not equal 4 therefore ordered pair (0, -1) is not on line 2x + 5y = 4

OpenStudy (lovelyanna):

Hmm.. so the answer does not have negative?

jigglypuff314 (jigglypuff314):

no... more like the answer cannot have x=0

OpenStudy (lovelyanna):

Oh.. so that leaves A, and B out, right?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Ok, give me a minute to think about it

OpenStudy (lovelyanna):

My head hurts.. :/ Can you explain more, please? :P

jigglypuff314 (jigglypuff314):

I'm use same example so for example for (0, -1) --> (x, y) so plug in 0 for x and -1 for y into 2x + 5y = 4 2x + 5y = 4 --> 2(0) + 5(-1) = 4 multiply and simplify 0 + -5 = 4 -5 does not equal 4 therefore ordered pair (0, -1) is not on line 2x + 5y = 4 another example: for (0, 1) plug in 0 for x and 1 for y 2x + 5y = 4 --> 2(0) + 5(1) = 4 simplify so 5 = 4 but 5 does not equal 4 therefore ordered pair (0, 1) is not one the line 2x + 5y = 4

OpenStudy (lovelyanna):

2x+5y=4. 2(7)=14 and 5(-2)=-10. 14+10=-34. -34 doesn't equal 4. Is it something like this, or did I get it wrong?

OpenStudy (lovelyanna):

14+(-10)= -34.* My bad

jigglypuff314 (jigglypuff314):

lol but then 14 - 10 = 4 right?

OpenStudy (lovelyanna):

Ooh.. right. :P Did I get everything right besides that? Then C

jigglypuff314 (jigglypuff314):

Correct ;)

OpenStudy (lovelyanna):

Thanks! :) Are you busy right now?

jigglypuff314 (jigglypuff314):

psh other than the fact that I have to finish all my homework in one hour because I have a concert to go to latter? of course not! xD If you need help, I can help though :)

OpenStudy (lovelyanna):

Lol, no, nvm. You do your homework! :P

OpenStudy (lovelyanna):

Btw, your very popular! :3 You have 91 fans!! (One of them is me! =3)

jigglypuff314 (jigglypuff314):

thanks ;) not as popular as some people though :P my favorite person is http://openstudy.com/users/hartnn <3 he's great!

OpenStudy (lovelyanna):

Cool! =) And Omg.. he has more than 1,000 fans.. :o *faints*

jigglypuff314 (jigglypuff314):

he saved my life on more that one occasion :P because I am currently failing my math class... I got a 52% on the first test, 57% on the second test, and a 44% on the quiz... >,<

OpenStudy (lovelyanna):

Failing? But your amazing in my math.. what grade are you in?

jigglypuff314 (jigglypuff314):

11th grade... APCalcAB is what I'm failing... started failing after trying to learn what is the slope of a curve... ... ...

OpenStudy (lovelyanna):

Oh, I do NOT know what it is, by the look of its name it looks hard.. :o Oh, soo sorry.. :( I wish I can help..

jigglypuff314 (jigglypuff314):

APCalcAB is "Advanced Placement Calculus" in other words, a college level course. :P I found that there are are general groups of people in the math section of this site ;) Those who are home schooled. Those who missed a class. Those who don't understand word problems. Those who need help with Calculus and Those who can help with all that xD

OpenStudy (lovelyanna):

Wow, must be hard, eh? :P Lol !! Gl with all that, just by reading them my head hurts! xD

OpenStudy (lovelyanna):

Btw, if you can, I want you to check your email everyday when your free, we can chat there instead of here! :)

jigglypuff314 (jigglypuff314):

accept invitation to chat? :P

OpenStudy (lovelyanna):

what invitation?

jigglypuff314 (jigglypuff314):

on gmail to chat? the chat bar...

OpenStudy (lovelyanna):

I dont know what that is, my friend and I usually chat like how I sent you an email, it gets full really fast! :P

jigglypuff314 (jigglypuff314):

okay then ;)

OpenStudy (lovelyanna):

Yay! :D So, whats this concert for/about?

jigglypuff314 (jigglypuff314):

I'm playing piccolo/flute in the concert :P

OpenStudy (lovelyanna):

Ooooh, cool =D I wanna here it?! D': Btw, you think I could see a pic of you without the hat covering half your face? :P

jigglypuff314 (jigglypuff314):

stalker much? lol jk I'll email you...

OpenStudy (lovelyanna):

Lol, naah :P Ok, I was just asking, you dont have to if you dont want to :P

OpenStudy (lovelyanna):

I need to go now.. Gl with your concert! :)

jigglypuff314 (jigglypuff314):

thanks :)

OpenStudy (lovelyanna):

Bye! :)